OpenStudy (anonymous):

If f(x) = 1/9x - 2, what is f^-1 (x)

OpenStudy (alekos):

still need help?

OpenStudy (clalgee):

rearrange equation to solve for x switch x and y to get the inverse function Let y = 1/(9x - 2) rearrange to solve for x 9x - 2 = 1/y 9x = 1/y + 2 x = 1/(9y) + 2/9 switch x and y y = 1/(9x) + 2/9 f⁻¹(x) = 1/(9x) + 2/9 ===== Let y = (1/9)x - 2. (1/9)x = y + 2 x = 9y + 18 g(x) = 9x + 18 <- Your final answer

OpenStudy (alekos):

how do you know that it's 1/(9x-2) ? Could be (1/9x) - 2 or (1/9)x - 2